OpenStudy (anonymous):

A sample of Geometry class final exam grades are as follows: 80, 85, 72, 76, 95 Find the variance. 1) 7.96 2) 49.23 3) 63.44 4) 6.72

OpenStudy (amistre64):

for variance we need the mean first

OpenStudy (anonymous):

what mean means

OpenStudy (amistre64):

\[var=\frac{\sum (x-\bar x)^2}{n-1}\]

OpenStudy (amistre64):

your asking a question that presupposes you know what a mean is; its the average of the given data

OpenStudy (anonymous):

add it all together then divide by 5

OpenStudy (amistre64):

yes, and that is what we use to get a final answer for variance.

OpenStudy (amistre64):

(mean - x)^2 = (x-mean)^2 so i see no need to be strict about that part of it tho

OpenStudy (anonymous):

its 332

OpenStudy (amistre64):

hmm, im not so sure about that, let me do some calculations in a general sense

OpenStudy (anonymous):

the answer is 49.23

OpenStudy (amistre64):

mean = (80+85+72+76+95)/5 = 81.6 (81.6-80)^2+(81.6-85)^2+(81.6-72)^2+(81.6-76)^2+(81.6-95)^2 = 317.2 317.2/(5-1) = 79.3 as a variance

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (amistre64):

then the answer choices are wrong.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%7B80%2C+85%2C+72%2C+76%2C+95%7D variance is the square of the sample standard deviation: (8.905)^2 = abt 76.29